OpenStudy (anonymous):

Tammy usually spends 10 minutes walking to school at a speed of 60 m/min. If she leaves home 2 min. Later than usual, then at what speed must she walk to reach school on time?

OpenStudy (anonymous):

okay, use the equation \[s=\frac{d}{t}\] we know the speed and time so find distance= which is 60 * 10=600m from home to school now she gets 2 min late, so her time spendng reduces (10-2)=8 min(which she has to manage to the school on time) we found the distance so 600m/8min= 75 m/min